OpenStudy (mckenzieandjesus):

check my answer please Find the domain of y=sqrt3x-8

OpenStudy (mckenzieandjesus):

\[x \ge \left(\begin{matrix}8 \\ 3\end{matrix}\right) <----- my answer\]

OpenStudy (mckenzieandjesus):

\[x \ge \left(\begin{matrix}3 \\ 8\end{matrix}\right)\]

OpenStudy (mckenzieandjesus):

\[x \ge -\left(\begin{matrix}8 \\ 3\end{matrix}\right)\]

OpenStudy (mckenzieandjesus):

\[x>-\left(\begin{matrix}8 \\ 3\end{matrix}\right)\]

zepdrix (zepdrix):

Stuff under the root needs to be `greater than or equal to 0`,\[\large\rm 3x-8\ge0\]And then you need to solve for x. You came up with positive 8 over 3? Hmm sounds right! :)

OpenStudy (mckenzieandjesus):

okay thanku